# Advances in Tabulating Carmichael Numbers

### v1

@@ Abstract.
We report that there are $49483968$ Carmichael numbers less than $10^{22}$ which is an order of magnitude improvement on the previous bound. The tabulation finds Carmichael numbers of the form $n=Pqr$ and is bifurcated by the size of $P$ with respect to the tabulation bound $B$. For $P<7\cdot 10^{7}$, we found $6680974$ Carmichael numbers and $985268$ of them were less than $10^{22}$. When $P>7\cdot 10^{7}$, we found $48498700$ Carmichael numbers. We provide a comprehensive overview of both cases of the algorithm. For the large case, we show and implement asymptotically faster ways to tabulate compared to the prior tabulation. We also provide an asymptotic estimate of the cost of this algorithm. The worst case $P$ are Carmichael numbers. A robust analysis of this algorithm would likely require resolution of long-standing open questions regarding their asymptotic density.
@@ 1. Introduction
Fermat’s Little Theorem states that when $p$ is prime that $a^{p}\equiv a\pmod{p}$ for any integer $a$. The converse of this theorem gives a computationally efficient way to detect if an integer is composite. Pick an integer $a<n$ and compute $a^{n}\pmod{n}$. If the result is not $a$, conclude that $n$ is composite. Unfortunately, there are composite integers for which this test will fail to detect as composite, e.g. $n=341$ and $a=2$. Even worse, there are composite number for which the congruence is $a$ for any $a<n$ and the least such example is $561$. It would be reasonable to call these numbers absolute Fermat pseudoprimes or S̆imerka numbers since V. S̆imerka found the first seven examples in 1885 [19]. However, we will keep with convention and call them Carmichael numbers.
@@
For a background and survey on Carmichael numbers we refer to [16]. While that survey is a bit dated, it shows the breadth of questions that have been asked. Our chief concern regards tabulation and the key source is R.G.E. Pinch’s The Carmichael numbers up to $10^{15}$ [10]. That paper also gives a background of prior tabulations of Carmichael numbers. In a series of reports, he announced further tabulations using the same algorithm with the most recent being up to $10^{21}$ [11, 12, 13, 14]. We address the problem of tabulation by investigating the difficulty of finding primes $q$ and $r$ given $P$ so that $Pqr$ is a Carmichael number.
@@
We analyze this problem from two different perspectives: bounded and unbounded. In the former case, we construct $q$ and $r$ so that $Pqr<B$. In the latter case, we analyse the cost to find any such primes so that $Pqr$ is a Carmichael number. The unbounded problem is interesting on its own and arises naturally in the context of a bounded computation. If $P$ is small enough with respect to $B$, any such $q,r$ that has $Pqr$ being a Carmichael number will inherently satisfy $Pqr<B$.
@@
Our contributions are as follows. We give a comprehensive overview and an analysis of tabulations algorithms. We review the improvements to the small case found in [18] and show a better way of handling the large case. In particular, there are specific cases in the tabulation which would have required an exponential amount of work but we give multiple ways to reduce this to only a polynomial amount of work (see Section 5). We also give a heuristic estimate on the amount of time to tabulate up to $B$ is
@@
See Theorem 8 in section 7 for a precise statement and argument.
@@
In [7], it is recounted how Shanks and Erdős disagreed regarding the asymptotic count of Carmichael numbers. Shanks had plausible beliefs informed by available empirical data. Erdős was informed by a heuristic arising from a theoretical construction technique. The conflict between empirical data and theory is reflected in this very paper. A practical algorithm for tabulation should be guided by empirically observation. When we make arguments based off of this data, we will speak of a guide or guidance. However, we are also informed by theory and do not want to give the impression that the guidance is a theorem (e.g. subsection 4.4). At the same time, we are concerned with asymptotic analysis of algorithms and offer asymptotic improvements whose benefits might not be empirically noticed in our current computational domain.
@@
$B\rm{exp}\left(\frac{-k\log B\log\log\log B}{\log\log B}\right)$
for some positive constant $k$. One interesting algorithmic consequence of accepting this viewpoint is that no tabulation can ever be in the “asymptotic domain”. Each successive tabulation would be more expensive than the last and would still be unrepresentative of what Carmichael numbers “truly” look like (e.g. they are always denser further out).
The Erdős construction starts with a carefully selected composite number $L$, and builds Carmichael numbers around $L$. There exist algorithmic versions of this construction technique and they have produced Carmichael numbers with billions of digits (e.g. [1]). A tabulation algorithm should be able to account for how such a number would be found in a tabulation, even if only in a hypothetical way. We show how these very numbers play a role as worst-case inputs to our tabulation algorithm (see section 5.5). Thus, we would likely need the asymptotic count of these numbers to prove a strong bound on the algorithmic cost and we find it especially pleasing to see this problem show up in an intrinsic way in the algorithm. If we were strictly guided by empirical observation it would have been more difficult to see this connection.
The remainder of the paper is organized as follows. In section 2, we review some general number-theoretic results that are needed and the specific results on Carmichael numbers. In section 3, we provide a high-level overview of the algorithm. In section 4, we review the “small” case and discuss its benefits. In section 5, we explain the “large” case. This includes Pinch’s approach and improvements based on divisors in residue classes[9, 2]. In Section 6, we discuss how construction of preproducts differs between the “small” and the “large” cases. In section 7, we provide a theorem on the count of preproducts. This serves as a guide for the total amount of work our tabulation algorithm does. Finally, we conclude with a report on our implementation and the tabulation to $B=10^{22}$. This independently verifies Pinch’s prior tabulations and exceeds his by an order of magnitude.
@@ Acknowledgements
We thank Richard Pinch, Carl Pomerance, and Jon Sorenson for helpful discussions that have made this paper better. We also thank Sungjin Kim for his help during a problem session at the 2022 West Coast Number Theory Conference. Thanks also to Frank Levinson for generous donations to Butler University that funded the BigDawg compute cluster. The second author is also grateful to the Holcomb awards committee for the financial support of this project.
@@ 2.1. Number Theoretic functions
This function shows up frequently in the tabulation algorithm. We are influenced by Theorem 2 and 3 of [4] and Theorem 5 of [6]. A guiding and colloquial summary of these results might be “$\lambda(n)$ is frequently close to $n$ in size, and $\lambda(n)$ may be very small infrequently.” We forgo precise statements because they are not used explicitly and we leave the reader to consult the sources. Even though we assume, as guidance, that $\lambda(n)$ is frequently close to $n$ in size, we design an algorithm that allows very small values of $\lambda(n)$. Indeed, we will see exactly how and where the small values play a significant role in the algorithm.
@@ 2.2. Theorems for Carmichael numbers
The tabulation methods construct $n$ in factored form. Our goal will be to construct square-free odd numbers $n$ which will then be tested with Korselt’s criterion. Let $d$ be the number of factors in $n$ and then $d>2$. Let $P_{k}=\prod_{i=1}^{k}p_{i}$ where $p_{i}<p_{j}$ iff $i<j$. The primary tabulation methods concern $P_{d-2}$ and constructing or searching for $p_{d-1}$ and $p_{d}$. Since these quantities are used so often, we write these quantities as $P$, $q$, and $r$.
@@
The immediately above theorem and corollary were the primary motivation of [18]. Given the finiteness result, it makes sense to tabulate with respect to $P$. Once a given $P$ is accounted for, it never needs to be revisited. The permanence of the tabulation may help reduce the work of future tabulations. As it stands, the below sections show that the above theorem is used in an algorithmic way when $P$ is accounted as “small” with respect to the total tabulation bound.
@@
Note that these bounds imply that a unbounded direct search is $\omega(P^{2})$ due to the cost of finding $q$. Trivially, the total cost is $O(P^{5})$. However, it is difficult to pin down the cost of finding $r$ given $P$ and $q$ due to the role $\lambda(P_{d-1})$ plays. We say more on this in section 5. From an implementation point of view, we will always have access to the value of $\lambda(P_{d-1})$ because $P_{d-1}$ is always constructed in factored form.
@@ 2.3. Computational Model
We are concerned with the counts of arithmetic operations used to generate triples $(P,q,r)$ and not concern ourselves with the asymptotic cost of verifying if $Pqr$ is a Carmichael number. The primary reason we omit the cost of verification is that the cost is not uniform across the computation. One could think of it as a claim that the tests are done in $O(1)$ time. The actual cost of the check varies significantly but we believe, on average, that the cost is $O(1)$. The primality test is usually the most expensive part of the check and may be deferred until other checks have passed. In section 4, we will generate $q$ and $r$ that are not integers and thus may be discarded with a single division. In section 5, we generate a prime $q$ by means of a look-up table and this is also a $O(1)$ cost. In all cases, we can check the $d$ divisibility statements associated with Korselt’s criterion before the primality tests. This uses at most $d$ divisions and $d<14$ for our particular computation. Even when all of these pass, and we need to verify the primality by use of a more expensive test, in some cases we know the complete (or partial) factorizations of $q-1$ or $r-1$ and have faster primality tests available.
We will account for storage costs in the bit model. And, as is typical in tabulation algorithms, we do not count the cost of storing the output.
@@ 3. Algorithmic Overview
while $P<X$ do $\triangleright$ See section 6.1.
@@
end while
while $P\geq X$ and $Pp_{d-2}^{2}<B$  do $\triangleright$ See section 6.2.
     for $q$ a prime in $(p_{d-2},\sqrt{B/P})$ do
@@
              Factor $Pq-1$. $\triangleright$ See sections 5.2-5.5
@@ 4. When $P$ is small
We present a concise and a clearer exposition of the key ideas found in [18]. In particular, the proofs of the run-times make it a bit more clear what the implied constants are for the computation associated with for a given $P$.
4.1. $CD$ method
The starting point is the third paragraph of section 2 of [10] where the “small” case is described. Theorem 3 is used directly to create the values $C$ and $D$ with a double-nested loop. The outer loop is on $D$ and the inner-loop on $C$ permitted by inequality $(3)$ of Theorem 3. With $C$, $D$, $\Delta$, the inner-loop does two integer checks (on $q$ and $r$), an optional bound check on $Pqr$, Korselt divisibility checks, and two primality checks (on $q$ and $r$).
In [18], it was shown that the number of $CD$ pairs created is akin to running a sieve of Eratosthenes on the interval found in the inequality $(3)$ of Theorem 3. Letting the length of that interval be $L_{P}$, then its value is about $2PP_{d-3}$. Note that when $d=3$, we interpret $P_{d-3}$ as the empty product and has value $1$. When $d=3$, $L_{p}$ is linear in $P$. As $d$ grows larger, $L_{P}$ can approach quadratic in $P$ because $P_{d-3}$ can approach $P^{1-\frac{1}{d-2}}$.
@@ 4.1. $CD$ method
The first equality is just rewriting $L_{P}$ in terms of $P$ and $P_{d-3}$ and would also be a $\Theta$ bound. The second equality is actually a containment statement of big-$O$ costs, since it represents an upper-bound on $P_{d-3}$. This method performs poorly when $P_{d-3}$ is large with respect to $P$.
4.2. $D\Delta$ method
The improvement found in [18], is to replace the loop on $C$ with a loop on $\Delta$, a constructed divisor of $(P-1)(P+D)$. The divisors are obtained using a variant of the sieve of Eratosthenes. The factorization of $(P-1)$ remains constant through all $D$ and the sieve variant factors all numbers in $[P+2,2P-1]$ which account for any $(P+D)$ term. While the cost of obtaining the divisors exceeds the cost to increment $C$, the fact that there are asymptotically fewer of the divisors makes up for increased overhead used to enter the inner loop. Much like the inner-loop of the $CD$-method, this approach does two integer checks (on $C=(P^{2}+\Delta)/D$ and $r$), an optional bound check, Korselt divisibility checks, and two primality checks (on $q$ and $r$).
@@ 4.2. $D\Delta$ method
$=$ $\tau(P-1)\left(P\log(P)\right)+O(P).$
There are a few practical improvements to the above to further reduce the number of divisors constructed. We may consider divisors of $(P-1)/2$ instead of $(P-1)$ because this forces $q$ to be an odd integer. The set of divisors may be further reduced considering a small prime $p$ dividing $D$ and only generating $\Delta$ satisfying $P^{2}\equiv-\Delta\pmod{p}$ for small primes $p$ dividing $D$. If $P\equiv 0,1\pmod{p}$, we either always include or exclude $p$ from the odometer generating the divisors. It is not too hard to see that both of these ideas only improve the constant term associated with the big-$O$ cost.
In [18], it was shown that the average cost, per $P$ is $O(P(\log P)^{2})$. We do not know how $\tau((P-1)/2)$ behaves, on average, compared to a “typical” integer since we are restricting the inputs to cyclic numbers.
4.3. Hybrid Method
Since the outer-loop on $D$ is the same in both methods, one may choose which inner-loop is the more optimal to enter. It is easy to compute the respective number of $CD$ or $D\Delta$ pairs that would have to be considered for a given $D$. Roughly speaking, the choice will be between the smaller of $2PP_{d-3}/D$ or $\tau((P-1)(P+D))$. While the former product decreases in a predictable way, the latter product can behave more chaotically. That said, it is closely related to $\tau(P-1)$ and since $P$ may be a Carmichael number this can take larger than expected values. Generally, when $D$ is small the $D\Delta$ method will be chosen. As $D$ gets larger, it may be that the $CD$ method becomes the faster way to complete the computation. As a guide, $P$ being prime is still common; so, we found that implementing this hybrid method was helpful in reducing actual run-time. In [18], it is conjectured that the hybrid method considers $O(P\log\log P)$ $CD$ or $D\Delta$ pairs when $P$ is a prime. If true, the hybrid method is an asymptotically faster algorithm for these cases.

### v2

@@ Abstract.
We report that there are $49679873$ Carmichael numbers less than $10^{22}$ which is an order of magnitude improvement on the previous bound. The tabulation finds Carmichael numbers of the form $n=Pqr$ and is bifurcated by the size of $P$ with respect to the tabulation bound $B$. For $P<7\cdot 10^{7}$, we found $35985331$ Carmichael numbers and $1202914$ of them were less than $10^{22}$. When $P>7\cdot 10^{7}$, we found $48476959$ Carmichael numbers less than $10^{22}$. We provide a comprehensive overview of both cases of the algorithm. For the large case, we show and implement asymptotically faster ways to tabulate compared to the prior tabulation. We also provide an asymptotic estimate of the cost of this algorithm. It is interesting that Carmichael numbers are worst case inputs to this algorithm. So, providing a more robust asymptotic analysis of the cost of the algorithm would likely require resolution of long-standing open questions regarding the asymptotic density of Carmichael numbers.
@@ 1. Introduction
Fermat’s Little Theorem states that when $p$ is prime that $a^{p}\equiv a\pmod{p}$ for any integer $a$. The converse of this theorem gives a computationally efficient way to detect if an integer is composite. Pick an integer $a<n$ and compute $a^{n}\pmod{n}$. If the result is not $a$, conclude that $n$ is composite. Unfortunately, there are composite integers for which this test will fail to detect as composite, e.g. $n=341$ and $a=2$. Even worse, there are composite numbers for which $a^{n}\equiv a\pmod{n}$ for any $a$, and the least such example is $561$. It would be reasonable to call these numbers absolute Fermat pseudoprimes or S̆imerka numbers since V. S̆imerka found the first seven examples in 1885 [26]. However, we will keep with convention and call them Carmichael numbers.
@@
For a background and survey on Carmichael numbers we refer to [21]. While that survey is a bit dated, it shows the breadth of questions that have been asked. Our concern is tabulating Carmichael numbers and this specific problem has a long pedigree [25, 22, 12, 13, 9]. The tabulation bounds for those sources were (respectively): $10^{9}$, $25\cdot 10^{9}$, $10^{12}$, $10^{13}$, and $10^{14}$. Our key source is R.G.E. Pinch’s The Carmichael numbers up to $10^{15}$ [15]. He also published a series of reports in which he announced further tabulations all using the same algorithm [16, 17, 18, 19]. The final tabulation found all Carmichael numbers less than $10^{21}$. We address the problem of tabulation by investigating the difficulty of finding primes $q$ and $r$ given $P$ so that $Pqr$ is a Carmichael number.
@@
We analyze this problem from two different perspectives: bounded and unbounded. In the former case, we construct $q$ and $r$ so that $Pqr<B$. In the latter case, we analyse the cost to find any such primes so that $Pqr$ is a Carmichael number. The unbounded problem is interesting on its own and arises naturally in the context of a bounded computation. If $P$ is small enough with respect to $B$, any primes $q$ and $r$ for which $Pqr$ is a Carmichael number will inherently satisfy $Pqr<B$.
@@
Our contributions are as follows. We give a comprehensive overview and an analysis of tabulation algorithms. The algorithm is bifurcated by size. For the small case, we review the improvements introduced in [24]. We do this for the sake of completeness and because the algorithms played a significant role in this tabulation. In [24], it was reported that the algorithm was run on $P<3\cdot 10^{6}$ and now this has been extended to $P<7\cdot 10^{7}$. In contrast, this paper shows a better way to handle the large case. In particular, there are specific cases in the tabulation which would have required an exponential amount of work but we give multiple ways to reduce this to only a polynomial amount of work (see Section 5). We also give a heuristic estimate on the amount of time to tabulate up to $B$, namely
$B\exp\left(\left(-\frac{1}{\sqrt{2}}+o(1)\right)\sqrt{\log B\log\log B}\right),$
@@
by providing an estimate on the number of preproducts the large case has to consider. See Theorem 8 in Section 7 for a precise statement and argument.
@@
In [8], it is recounted how Shanks and Erdős disagreed regarding the asymptotic count of Carmichael numbers. Shanks had plausible beliefs informed by available empirical data. Erdős was informed by a heuristic arising from a theoretical construction technique. The conflict between empirical data and theory is reflected in this very paper. A practical algorithm for tabulation should be guided by empirically observation. When we make arguments based off of this data, we will speak of a guide or guidance. However, we are also informed by theory and do not want to give the impression that the guidance is a theorem (e.g., Subsection 4.4). At the same time, we are concerned with asymptotic analysis of algorithms and offer asymptotic improvements whose benefits might not be empirically noticed in our current computational domain.
@@
$B\exp\left(\frac{-k\log B\log\log\log B}{\log\log B}\right)$
for some positive constant $k$. This implies that the count is greater than $B^{1-\epsilon}$ for any fixed $\epsilon>0$. When Alfords, Pomerance, and Granville proved the infinitude of Carmichael numbers, they also established a lower bound $B^{2/7}$ for their count[2]. The current best lower bound is $B^{1/3}$[10], which is more empirically in line with what current tabulations show. We seem quite far from the point at which tabulations could give empirical evidence for Erdős’ heuristic. This dichotomy was also explicitly pointed out by Swift in 1975 [25].
The Erdős construction starts with a carefully selected composite number $L$, and builds Carmichael numbers around $L$. There exist algorithmic versions of this construction technique that produced Carmichael numbers with billions of prime factors (e.g., [1]). A tabulation algorithm should be able to account for how such a number would be found, even if only in a hypothetical way. We show the sorts of numbers that arise from this construction technique play a role as worst-case inputs to our tabulation algorithm (see Subsection 5.5). Thus, we would likely need the asymptotic count of these numbers to prove a strong bound on the algorithmic cost and we find it especially pleasing to see this problem show up in an intrinsic way in the algorithm. If we were strictly guided by empirical observation it would have been more difficult to see this connection.
The remainder of the paper is organized as follows. In Section 2, we review some general number-theoretic results that are needed and the specific results on Carmichael numbers. In Section 3, we provide a high-level overview of the algorithm. In Section 4, we review the small case and discuss its benefits. In Section 5, we explain the large case. This includes Pinch’s approach and improvements based on divisors in residue classes [14, 3]. In Section 6, we discuss how construction of preproducts differs between the small and the large cases. In Section 7, we provide a theorem on the count of preproducts. This serves as a guide for the total amount of work our tabulation algorithm does. Finally, we conclude with a report on our implementation and the tabulation to $B=10^{22}$. This independently verifies Pinch’s prior tabulations and exceeds his by an order of magnitude.
@@ Acknowledgements
We thank Richard Pinch, Carl Pomerance, and Jon Sorenson for helpful discussions that have made this paper better. We thank Drew Sutherland for his generosity and the Simons Collaboration in Arithmetic Geometry, Number Theory, and Computation for computational support. We also thank Sungjin Kim for his help during a problem session at the 2022 West Coast Number Theory Conference. Thanks also to Frank Levinson for generous donations to Butler University that funded the BigDawg compute cluster. The second author is also grateful to the Holcomb Awards Committee for the financial support of this project.
@@ 2.1. Number Theoretic functions
This function shows up frequently in the tabulation algorithm. We are influenced by Theorem 2 and 3 of [5] and Theorem 5 of [7]. A guiding and colloquial summary of these results might be, “$\lambda(n)$ is frequently close to $n$ in size, and $\lambda(n)$ may be very small infrequently.” We forgo precise statements because they are not used explicitly and we leave the reader to consult the sources. Even though we assume, as guidance, that $\lambda(n)$ is frequently close to $n$ in size, we design an algorithm that allows very small values of $\lambda(n)$. Indeed, we will see exactly how and where the small values play a significant role in the algorithm.
@@ 2.2. Theorems for Carmichael numbers
The tabulation methods construct $n$ in factored form. Our goal will be to construct square-free odd numbers $n$ which will then be tested with Korselt’s criterion. Letting $d$ be the number of prime factors dividing $n$, then $d>2$. Let $P_{k}=\prod_{i=1}^{k}p_{i}$ where $p_{i}<p_{j}$ iff $i<j$. The primary tabulation methods concern $P_{d-2}$ and constructing or searching for $p_{d-1}$ and $p_{d}$. Since these quantities are used so often, we write these quantities as $P$, $q$, and $r$. For reference, this means that $P_{d-1}=Pq$ and $P_{d}=Pqr$.
@@
The immediately above theorem and corollary were the primary motivation of [24]. Given the finiteness result, it makes sense to tabulate with respect to $P$. Once a given $P$ is accounted for, it never needs to be revisited. The permanence of the tabulation may help reduce the work of future tabulations. As it stands, the below sections show that the above theorem is used in an algorithmic way when $P$ is accounted as small with respect to the total tabulation bound.
@@
The computational cost of an unbounded direct search is $\Omega(P^{2})$ due to the cost of finding $q$. It is difficult to pin down the cost of finding $r$ given $P$ and $q$ due to the role $\lambda(P_{d-1})$ plays. We say more in Section 5. An implementation will always have access to $\lambda(P_{d-1})$ because $P_{d-1}$ is always constructed in factored form.
@@ 2.3. Computational Model
We are concerned with the counts of arithmetic operations used to generate triples $(P,q,r)$ and do not concern ourselves with the asymptotic cost of verifying if $Pqr$ is a Carmichael number. The primary reason we omit the cost of verification is that the cost is not uniform across the computation. We will treat the cost of verification as being done in constant time as a guide.
The primality test is the most expensive part of the Korselt check and we deferred it until other checks have passed. This includes integrality checks for $q$ and $r$ and the $d<14$ divisibility statements in the Korselt criterion. We may also discard candidates beyond the tabulation bound. Even when all of these pass, and we need to verify the primality by use of a more expensive test, in some cases we know the complete (or partial) factorizations of $q-1$ or $r-1$ and have faster primality tests available.
We will account for storage costs using the bit model. And, as is typical in tabulation algorithms, we do not count the cost of storing the output.
@@ 3. Algorithmic Overview
for each $P<X$ do $\triangleright$ See Section 6.1.
@@
end for
for each $P$ satisfying $P\geq X$ and $Pp_{d-2}^{2}<B$  do $\triangleright$ See Section 6.2.
     for each $q$ an admissible prime in $(p_{d-2},\sqrt{B/P})$ do
@@
              Factor $Pq-1$ using divisors in residue classes. $\triangleright$ See Section 5.3
@@ 4. When $P$ is small
We present a concise and clear exposition of the key ideas found in [24]. In particular, the proofs of the run-times make it a bit more clear what the implied constants are for the computation associated for a given $P$.
4.1. The $CD$ method
The starting point is the third paragraph of Section 2 of [15] where the small case is described. Theorem 3 is used to create the values $C$ and $D$ with a double nested loop. The outer loop is on $D$ and the inner loop is on $C$ permitted by inequality $(3)$ of Theorem 3. With $C$, $D$, and $\Delta$, the inner loop does divisibility checks on $q$ and $r$, an optional bound check on $Pqr$, Korselt divisibility checks, and two primality checks.
In [24], it was shown that the number of $CD$ pairs created is akin to running a sieve of Eratosthenes on the interval found in the inequality $(3)$ of Theorem 3. Letting the length of that interval be $L_{P}$, then its value is about $2PP_{d-3}$. Considering the $d=3$ case, we have $P_{d-3}=1$ and so, $L_{p}$ is linear in $P$. As $d$ grows larger, $L_{P}$ can approach quadratic in $P$ because $P_{d-3}$ can approach $P^{1-\frac{1}{d-2}}$.
@@ 4.1. The $CD$ method
The first equality is just rewriting $L_{P}$ in terms of $P$ and $P_{d-3}$, and would also be a $\Theta$ bound. The second equality is a containment statement of big-$O$ costs, since it represents an upper bound on $P_{d-3}$. This method performs poorly when $P_{d-3}$ is large with respect to $P$.
4.2. The $D\Delta$ method
The improvement found in [24], is to replace the loop on $C$ with a loop on $\Delta$, constructed divisors of $(P-1)(P+D)$. The divisors are obtained using a variant of the sieve of Eratosthenes. The factorization of $(P-1)$ remains constant through all $D$ and the sieve variant factors all numbers in $[P+2,2P-1]$, which accounts for the factors of any $(P+D)$ term. Even though the cost of obtaining the divisors exceeds the cost to increment $C$, the asymptotically fewer number of the divisors makes up for increased overhead used to enter the inner loop. Like the inner loop of the $CD$ method, this approach does divisibility checks on $C=(P^{2}+\Delta)/D$ and $r$, an optional bound check, Korselt divisibility checks, and two primality checks.
@@ 4.2. The $D\Delta$ method
$=$ $\tau(P-1)\left(P\log(P)+O(P)\right).$
There are a few practical improvements to the above to further reduce the number of divisors constructed. We may consider divisors of $(P-1)/2$ instead of $(P-1)$ because this forces $q$ to be an odd integer. The set of divisors may be further reduced considering a small prime $p$ dividing $D$ and only generating $\Delta$ satisfying $P^{2}\equiv-\Delta\pmod{p}$ for small primes $p$ dividing $D$. If $P\equiv 0,1\pmod{p}$, we either always include or exclude $p$ from the odometer generating the divisors. It is not too hard to see that both of these ideas only improve the implied constant on the big-$O$ cost.
4.3. The Hybrid Method
Since the outer loop on $D$ is the same in both methods, one may choose which inner loop is better to enter. It is easy to compute the number of $C$ values or the number of divisors $\Delta$ that would have to be considered for a given $D$. Roughly speaking, the choice will be between the smaller of $2PP_{d-3}/D$ or $\tau((P-1)(P+D))$. While $2PP_{d-3}/D$ decreases in a predictable way, the divisor count can behave more chaotically. Generally, when $D$ is small the $D\Delta$ method will be chosen. As $D$ gets larger, it may be that the $CD$ method becomes the faster way to complete the computation. It was common for $P$ to be prime in our computational range; so, we found that implementing this hybrid method was helpful in reducing run-time. In [24], it is conjectured that the hybrid method considers $O(P\log\log P)$ $CD$ or $D\Delta$ pairs when $P$ is a prime. If true, the hybrid method is an asymptotically faster algorithm for these cases.